OpenStudy (anonymous):

Find the solution to the equation log16 32 = x + 3

Directrix (directrix):

Log base 16 if 32 = (x + 3) means, by definition of logarithms that: 16 ^(x+3) = 32

Directrix (directrix):

Rewrite 16 as a power of 2. Do the same for 32.

Directrix (directrix):

2^4 = 16 2^5 = 32 The purpose of rewriting 16 and 32 as powers of 2 is to apply exponential properties in the solution process.

Directrix (directrix):

16 ^(x+3) = 32 (2^4)^(x+3) = 2^5 2^4(x+3) = 2^5 So, 4(x+3) = 5 Solve 4(x+3) = 5 and check to see if the resulting value of x works in the original equation. I will draw most of the solution by hand so that it will make more sense.

Directrix (directrix):